It is desired to obtain a volume of 1.00 L of nitrogen (N2) at 100.00 C and 640.00 mmHg. How many moles of nitrogen (N2) would be required


1
Expert's answer
2022-05-18T06:19:06-0400

As we know p1v1=p2v2, v2=p1v1/p2; 1*640/760= 0.842 liters, n=v/vm=0.842/22.4 = 0.0376 moles of N2 would be requried